# Towards Human-Level 3D Relative Pose Estimation: Generalizable, Training-Free, with Single Reference

### v1

@@ III-B Semantic Map Estimation
In order to estimate the relative pose, human intelligence may unconsciously infer the semantics of the reference-query pair. Subsequently, coarse correspondence can be established with those semantics, resulting in three-fold benefits: it (i) helps to filter out the large non-overlapped part under a substantial pose difference, (ii) alleviates the influence of occlusions, and (iii) eases the degraded optimization of the relative pose estimation.
Benefit from the rapid development of large pretrained models, an elegant off-the-shelf semantic feature extractor is available as DINO/DINOv2 dino ; dinov2 , which shows great zero-shot generalize-ability to diverse (even texture-less) objects (see Fig. 3 for some examples). We thus incorporate the off-the-shelf DINOv2 model dinov2 to acquire the rich semantics of the input unseen objects.
@@
By feeding Eq. (1) with $I_{q}$ and $I_{r}$, we can obtain the semantic maps for the query and the reference, $S_{q}$ and $S_{r}$, respectively.
@@
Specifically, given the depth map $D_{r}$ of the reference, we lift the coordinates of the image plane into the 3D space and obtain the 2.5D point clouds $X_{r}\in\mathbb{R}^{N\times 3}$ of the front surface. We then reconstruct the corresponding 2.5D mesh $M_{r}$ from $X_{r}$, to facilitate the rasterization in the renderer. Since the $xy$ coordinates of $X_{r}$ are sampled regularly from the 2D grids, reconstructing $M_{r}$ from $X_{r}$ can be easily achieved by the Delaunay triangulations lee1980two . Finally, we texture $M_{r}$ with both color and semantic maps, obtaining $\mathcal{M}_{r}=\texttt{TextMap}(M_{r},I_{r},S_{r})$ for rendering under novel poses.
Note that as discussed in Sect. I-A (Applicability), our method possesses the potential of using only an RGB reference and estimating an imprecise depth map exploiting an off-the-shelf generalizable depth estimator. Good estimation is validated in supplementary material given an imprecise and noising depth. We leave training a generalizable depth estimator as our future work to avoid possible distractions in this paper.
@@ III-D Label/Training-Free Refinement via Differentiable Renderer
Formally, by assuming a perspective camera, we leverage a recent differentiable renderer nvidiffrast nvidiffrast , denoted as $\mathcal{R}$, to generate novel RGB and semantic maps, $I_{r}(P)$ and $S_{r}(P)$, from the textured 2.5D reference mesh $\mathcal{M}_{r}$, an arbitrary pose $P$, and the camera intrinsics $K$:
@@
Initialization. As revealed in the majority of prior arts deepim ; megapose ; templates-pose ; latentfusion , a good initialization significantly boosts the performance of the render-and-compare framework. To this end, we implement our initialization by evenly sampling candidate poses on a sphere and chasing the best one. Specifically, we first sample $m$ viewpoints (azimuth and elevation angles) uniformly using a Fibonacci lattice gonzalez2010measurement , then uniformly sample $n$ in-plane rotation angles for each viewpoint, producing $t=m*n$ poses as the initializing candidates. By rendering both RGB and semantic maps of those candidate poses, we are able to calculate the re-projection loss by Eq. (3) (without back-propagation in this phase) and choose the pose with the minimal loss as our initialization $P^{\text{init}}$.
@@
0:  Reference RGB and depth $I_{r},D_{r}$; query RGB $I_{q}$; differentiable renderer $\mathcal{R}$; pretrained DINOv2 model $\mathbf{\Phi}$, iteration quota $N$, learning rate $\alpha$, camera intrinsic $K$.
1:  $S_{q}\leftarrow\texttt{PCA}(\mathbf{\Phi}(I_{q}))$, $S_{r}\leftarrow\texttt{PCA}(\mathbf{\Phi}(I_{r}))$
2:  $M_{r}\leftarrow\texttt{DelaunayTriangulations}(I_{r},D_{r})$
3:  $\mathcal{M}_{r}\leftarrow\texttt{TextMap}(M_{r},I_{r},S_{r})$triangleright Sampling Poses for Initialization
@@
5:  $\mathbf{P}=\left\{P^{\text{1}},P^{\text{2}},...,P^{n}\right\}$
6:  $I_{r}(\mathbf{P}),S_{r}(\mathbf{P})\leftarrow\mathcal{R}(\mathbf{P},\mathcal{M_{\text{r}}},K)$
@@ IV-A Experimental Setups
Evaluation Metric. Following zhao2024dvmnet and 3DAHV , we report mean angular error across sampled reference-query pairs. We also evaluate on important metrics of Acc@5/10/15/30, i.e., the percentage of the predictions that are within 5/10/15/30, which can be more rigorous (e.g., Acc@5) and better characterize the performance. The degree of the pose difference between the ground truth $R_{gt}$ and the predictions $\hat{R}$ is calculated by the geodesic distance $D$:
@@
Since the results on the rigorous metrics of Acc@5/10 are not reported in the 3DAHV 3DAHV and DVMNet zhao2024dvmnet paper, we thus retrain them using the code released by the authors for the Acc@5/10 evaluation.
Moreover, as a label/training-free method, the performance of our method can be assessed on all the objects of LineMOD, YCB-V, and LM-O datasets, without the need to leave out any training data or leverage any external dataset. We report the performance of our method on the complete LineMOD, YCB-V, and LM-O datasets in Tables S3, S4, and S5 of the supplementary material.
In-dataset and Cross-dataset Evaluation. Beyond the unseen objects generalization, we also test the dataset-level generalization for the state-of-the-art network-based methods DVMNet and 3DAHV, reporting both the in-dataset and the cross-dataset performance. In short, in-dataset and cross-dataset differ in whether the network needs to be finetuned on a subset that comes from the same dataset with the testing set (though not including the testing objects). Therefore, a good cross-dataset performance demonstrates better generalization in terms of the dataset, as the network only needs to be trained once on a large-scale external dataset without finetuning. Specifically, for the in-dataset experiments, we follow the exact training protocols of DVMNet zhao2024dvmnet and 3DAHV 3DAHV , which first pretrain on an external large-scale dataset Objaverse objaverse then fintune on a certain dataset (e.g., LineMOD or YCB-V). For cross-dataset experiments, we use the pretrained weights from Objaverse directly without finetuning.
Note that the evaluation of our method, ZSP, and LoFTR does not involve a finetuning phase, demonstrating that our method, ZSP, and LoFTR naturally generalize to an arbitrary dataset [^4] This is achieved by that 1) the pose estimation phase of our method, ZSP, and LoFTR are general and do not involve learning a network, and 2) they all use generalizable feature extractors, i.e., DINOv2 or LoFTR..
@@
Following DVMNet zhao2024dvmnet and 3DAHV 3DAHV , for each object, we generate 1000 pairs for testing, and 20000 pairs for training DVMNet, 3DAHV, and RelPose++. Figure 4 illustrates the histograms depicting the statistics of the pairwise pose difference (geodesic distance between rotation matrices $R_{r}$ and $R_{q}$) on the three datasets. All the experiments are carried out on the same testing reference-query pairs.
Implementation Details. For semantic feature extraction, we employ the output tokens from the last layer of the DINOv2 ViT-L model dinov2 . We use nvdiffrast nvidiffrast as our differentiable renderer. We uniformly sample $m=200$ viewpoints and $n=20$ in-plane rotations (resulting in 4000 initialization candidates), the maximal iteration number for differentiable rendering is set to $N=30$. To backpropagate the refinement losses, we use an Adam optimizer adam of 0.01 initial learning rate and decay by a ReduceLROnPlateau scheduler. All the experiments are conducted on a single NVIDIA 4090 GPU.
IV-B Experimental Results on the LineMOD Dataset
The results on the LineMOD dataset are illustrated in Table II. We paste the performances of RelPose++ from the 3DAHV paper 3DAHV . We leave the Acc@5/10 performance of RelPose++ blank as those were not reported in 3DAHV and the (pre-) training code of RelPose++ on the external large-scale Objaverse dataset is not available.
Table II shows that our label and training-free method significantly outperforms the supervised state-of-the-art DVMNet w.r.t. all the metrics. In addition, the state-of-the-art methods DVMNet and 3DAHV face challenges in generalizing across different datasets, i.e., their in-dataset results substantially outperform their cross-dataset counterparts. In contrast, our approach, without training a network, inherently generalizes across diverse datasets directly. Especially, our method significantly outperforms DVMNet (in-dataset) for 21.6% and 32.08% w.r.t. the rigorous Acc@5/10. The qualitative results of our method are shown in Fig. 5, and comparisons with different methods are presented in Fig. S3 of the supplementary material. Our results on all the LineMOD objects are detailed in Table S3 of the supplementary material.
TABLE II: Experimental results on LineMOD. We illustrate both the experimental settings and the performance. In the RGB-D category, both means requiring RGB-D image for both query and reference. Acc@$t^{\circ}$ measures the percentage of the estimated pose within $t^{\circ}$ w.r.t. the ground-truth.
@@ IV-D Experimental Results on the LM-O Dataset
We observe that our results in terms of Mean Err are inferior to the in-dataset results of the state-of-the-art DVMNet and 3DAHV (though our method exhibits better Acc@$t^{\circ}$ results). This can be attributed to the extensive occlusions presented in the LM-O dataset, which lead to numerous testing pairs lacking adequate overlap. Consequently, those testing pairs are difficult to handle by all the methods (and also challenging for humans). We show those samples as failure cases in Fig. 8 of Sect. V-E, as well as investigating the angle error distribution (ranging from 0 to 180 degrees) on the LM-O dataset in Fig. S1 of the supplementary materials. The statistics reveal that at lower angle error thresholds (e.g., for $t\leq 10,20$ in Acc@$t^{\circ}$), our approach substantially outperforms both DVMNet and 3DAHV. This indicates that for test pairs with sufficient overlaps (i.e., match-able testing pairs), our method delivers superior performance compared to the state-of-the-art DVMNet and 3DAHV.
@@
Method Settings Error↓ Acc @$t^{\circ}$ (%) ↑
@@ V Ablation Analysis
We carefully investigate the following issues by ablation: 1) the contribution of each comprising element of our method, including the back-surface culling, and the usage of RGB or semantic modality in Sect. V-A; 2) the effects of different initialization strategies in Sect. V-B; 3) the effects of different refinement iterations in Sect. V-C; 4) the inference time statistics of our method in Sect. V-D; and 5) the failure cases illustrations from the LM-O dataset in Sect. V-E.
V-A The Contributions of the Proposed Comprising Elements
As expected, removing each of our comprising elements results in a decreased performance, because all of them are exploited with clear motivations. Nonetheless, the encouraging observation is that our method is able to deliver promising results using only the RGB modality without the semantic map. This further extends the applicability of our method when the pretrained DINOv2 model is not available or when the DINOv2 model cannot produce reasonable outputs (though the latter case could be rare).
@@ V-A The Contributions of the Proposed Comprising Elements
Metrics Mean Err↓ Acc @$30^{\circ}$ Acc @$15^{\circ}$ Acc @$10^{\circ}$ Acc @$5^{\circ}$
@@
V-B Effects of Different Initialization Strategies
The pose estimation performance under the render-and-compare paradigm is largely affected by the initialization deepim ; cir ; rnnpose ; megapose ; templates-pose ; latentfusion . In the following, we investigate different initializations including: 1) random initialization, where we randomly sample candidate poses and choose the best one; and 2) uniform initialization, where the candidate poses are uniformly sampled from a Fibonacci lattice with in-plane rotations gonzalez2010measurement , as detailed in Sect. III-D (Initialization). For the latter, we also examine different densities of the sampling, i.e., the Fibonacci lattice viewpoints including 100 and 200, and in-plane rotations including 20 and 50.
Table VI illustrates the performance of different initialization strategies using the LineMOD dataset, which demonstrates that 1) the uniform initialization outperforms the random initialization, and 2) uniform initialization with denser sampling leads to better performance. In our experiments, we choose uniform initialization with 4000 samples (200 Fibonacci lattice viewpoints times 20 in-plane rotations) to balance the performance and the efficiency.
TABLE VI: Effects of different initialization strategies using the LineMOD dataset. We evaluate different initialization strategies (rows 1-2) as well as the sampling density (rows 3-5).
Inital Strategy Sampling Numbers Error↓ Acc @ $t^{\circ}$ (%) ↑
@@ V-B Effects of Different Initialization Strategies
Uniform Init. (used in our experiments) 4000 (200 viewpoints, 20 in-plane rotations) 29.93 72.06 54.90 42.74 24.32
Uniform Init. 2000 (100 viewpoints, 20 in-plane rotations) 32.51 69.36 49.26 38.80 21.70
Uniform Init. (used in our experiments) 4000 (200 viewpoints, 20 in-plane rotations) 29.93 72.06 54.90 42.74 24.32
Uniform Init. 10000 (200 viewpoints, 50 in-plane rotations) 29.13 73.30 57.46 44.92 25.02
V-C Effects of Different Refinement Iterations.
Table VII illustrates the impact of the iteration numbers for our label/training-free refinement using the LineMOD dataset. It shows that the improvement becomes marginal after the iteration number $N$ exceeds 30. We thus set the iteration number to $N=30$ to achieve a balance between performance and efficiency.
TABLE VII: Effects of different refinement iterations on LineMOD.
Iterations N Mean Err↓ Acc @$30^{\circ}$ Acc @$15^{\circ}$ Acc @$10^{\circ}$ Acc @$5^{\circ}$
10 30.46 71.92 53.40 39.98 19.86
20 30.16 72.04 54.18 41.84 22.60
30 (used in our exp.) 29.93 72.06 54.90 42.74 24.32
50 29.86 72.10 54.94 42.92 25.18
100 29.73 72.12 55.26 43.60 25.64
V-D The Statistics of Our Inference Time
We collect the inference time per reference-query pair, averaged across the LineMOD datasets on a single 4090 GPU. We report the runtime for each stage of our method in Table VIII. Note that the initialization is efficient with much more candidate samples than the refinement, because those initializing candidate samples can be evaluated in parallel without backpropagation. Table VIII demonstrates the efficiency of our method with a per-pair runtime of 4.85 seconds in total.
@@ V-D The Statistics of Our Inference Time
0.24 s 3.46 s 1.03 s 4.85 s
V-E Illustrations of the Failure Cases

### v2

@@ III-B Semantic Map Estimation
In order to estimate the relative pose, human intelligence may unconsciously infer the semantics of the reference-query pair. Subsequently, coarse correspondence can be established with those semantics, resulting in three-fold benefits: i) it helps to filter out the large non-overlapped part under a substantial pose difference, ii) alleviates the influence of occlusions, and iii) eases the degraded optimization of the relative pose estimation.
Benefit from the rapid development of large pretrained models, an elegant off-the-shelf semantic feature extractor is available as DINO/DINOv2 dino ; dinov2 , which shows great zero-shot generalizability to diverse (even texture-less) objects (see Fig. 3 for some examples). We thus incorporate the off-the-shelf DINOv2 model dinov2 to acquire the rich semantics of the input unseen objects.
@@
By feeding Eq. (1) with $I_{q}$ and $I_{r}$, we can obtain the semantic maps for the query and the reference, $S_{q}$ and $S_{r}$, respectively. To ensure the semantic consistency between reference and query images, we first calculate and fix the PCA transformation using the reference image, then apply this transformation to all the query images.
@@
Specifically, given the depth map $D_{r}$ of the reference, we lift the coordinates of the image plane into the 3D space and obtain the front surface 2.5D point clouds $X_{r}\in\mathbb{R}^{N\times 3}$. We then reconstruct the corresponding 2.5D mesh $M_{r}$ from $X_{r}$, to facilitate the rasterization in the renderer. Since the $xy$ coordinates of $X_{r}$ are sampled regularly from the 2D grids, reconstructing $M_{r}$ from $X_{r}$ can be easily achieved by the Delaunay triangulations lee1980two . Finally, we texture $M_{r}$ with both color and semantic maps, obtaining $\mathcal{M}_{r}=\texttt{TextMap}(M_{r},I_{r},S_{r})$ for rendering under novel poses.
Note that as discussed in Sect. I-A (Applicability), our method possesses the potential of using only an RGB reference and estimating an imprecise depth map exploiting an off-the-shelf generalizable depth estimator. Good estimation is validated in Sect. S2 of the supplementary material given an imprecise and noisy depth. We leave training a generalizable depth estimator as our future work to avoid possible distractions in this paper.
@@ III-D Label/Training-Free Refinement via Differentiable Renderer
Formally, by assuming a perspective camera, we leverage a recent differentiable renderer nvdiffrast nvdiffrast , denoted as $\mathcal{R}$, to generate novel RGB and semantic maps, $I_{r}(P)$ and $S_{r}(P)$, from the textured 2.5D reference mesh $\mathcal{M}_{r}$, an arbitrary pose $P$, and the camera intrinsic KK [^4] The camera intrinsic $K$ can be obtained from the image EXIF information.:
@@
Initialization. As revealed in the majority of prior arts deepim ; megapose ; templates-pose ; latentfusion , a good initialization significantly boosts the performance of the render-and-compare framework.
To this end, we implement our initialization by evenly sampling candidate poses on a sphere and chasing the best one. Specifically, we first sample $m$ viewpoints (azimuth and elevation angles) uniformly using a Fibonacci lattice gonzalez2010measurement , then uniformly sample $n$ in-plane rotation angles for each viewpoint, producing t=mnt=m*n poses as the initializing candidates. By rendering both RGB and semantic maps of those candidate poses, we are able to calculate the re-projection loss by Eq. (3) (without back-propagation in this phase) and choose the pose with the minimal loss as our initialization $P^{\text{init}}$.
@@
0: Reference RGB and depth $I_{r},D_{r}$; query RGB $I_{q}$; differentiable renderer $\mathcal{R}$; pretrained DINOv2 model $\mathbf{\Phi}$, iteration quota NN, learning rate $\alpha$, camera intrinsic $K$.
@@
[…]
@@ IV-A Experimental Setups
Evaluation Metric. Following zhao2024dvmnet ; 3DAHV , we report mean angular error across sampled reference-query pairs. We also evaluate on important metrics of Acc@5/10/15/30, i.e., the percentage of the predictions that are within 5/10/15/30, which can be more rigorous (e.g., Acc@5) and better characterize the performance. The degree of the pose difference between the ground truth $R_{gt}$ and the predictions $\hat{R}$ is calculated by the geodesic distance $D$:
@@
Since the results on the rigorous metrics of Acc@5/10 are not reported in the 3DAHV 3DAHV and DVMNet zhao2024dvmnet paper, we thus retrain them using their official codes for the Acc@5/10 evaluation.
Moreover, as a label/training-free method, the performance of our method can be assessed on all the objects of LineMOD, YCB-V, and LM-O datasets, without the need to leave out any training data or leverage any external dataset. We report the performance of our method on the complete LineMOD, YCB-V, and LM-O datasets in Tables S5, S6, and S7 of the supplementary material.
In-dataset and Cross-dataset Evaluation. Beyond the unseen objects generalization, we also test the dataset-level generalization for the network-based methods including the state-of-the-art DVMNet zhao2024dvmnet and 3DAHV 3DAHV , reporting both the in-dataset and the cross-dataset performance. In short, in-dataset and cross-dataset differ in whether the network needs to be finetuned on a subset that comes from the same dataset with the testing set (though not including the testing objects). Therefore, a good cross-dataset performance is desirable, as the network only needs to be (pre-) trained once on a large-scale external dataset without finetuning.
Specifically, for the in-dataset experiments, we follow the exact training protocols of DVMNet zhao2024dvmnet and 3DAHV 3DAHV , which first pretrain on an external large-scale dataset Objaverse objaverse then finetune on a certain dataset (e.g., LineMOD or YCB-V). For cross-dataset experiments, we use the pretrained weights from Objaverse directly without finetuning.
Note that our method, ZSP zsp , and LoFTR loftr do not require a finetuning phase before evaluation, suggesting that our method, ZSP, and LoFTR naturally generalize to an arbitrary dataset [^5] This is achieved by that i) the pose estimation phase of our method, ZSP, and LoFTR are general and do not involve learning a network, and ii) they all use generalizable feature extractors, i.e., DINOv2 or LoFTR.
@@
Following DVMNet zhao2024dvmnet and 3DAHV 3DAHV , for each object, we generate 1000 pairs for testing, and 20000 pairs for training DVMNet, 3DAHV, and RelPose++. Fig. 4 illustrates the histograms depicting the statistics of the pairwise pose difference (geodesic distance between rotation matrices $R_{r}$ and $R_{q}$) on the three datasets. All the experiments are carried out on the same testing reference-query pairs.
Implementation Details. For semantic feature extraction, we employ the output tokens from the last layer of the DINOv2 ViT-L model dinov2 . We use nvdiffrast nvdiffrast as our differentiable renderer. We uniformly sample $m=200$ viewpoints and $n=20$ in-plane rotations (resulting in 4000 initialization candidates), the maximal iteration number for differentiable rendering is set to $N=30$. To backpropagate the refinement losses, we use an Adam optimizer adam of 0.01 initial learning rate and decay by a ReduceLROnPlateau scheduler. All the experiments are conducted on a single NVIDIA 4090 GPU.
[…]
@@
IV-B Experimental Results on the LineMOD Dataset
The results on the LineMOD dataset are illustrated in Table II. We paste the performances of RelPose++ from the 3DAHV paper 3DAHV . We leave the Acc@5/10 performance of RelPose++ blank as those were not reported in 3DAHV and the (pre-) training code of RelPose++ on the external large-scale Objaverse dataset is not available. Table II shows that our label and training-free method significantly outperforms the supervised state-of-the-art DVMNet w.r.t. all the metrics. In addition, the experimental results show that the performance of DVMNet and 3DAHV decreases when facing cross-dataset scenarios. In contrast, without training a network, our approach inherently generalizes across diverse datasets. Especially, our method significantly outperforms DVMNet (in-dataset) for 21.6% and 32.08% w.r.t. the rigorous Acc@5/10. The qualitative results of our method are shown in Fig. 5, and comparisons with different methods are presented in Fig. S3 of the supplementary material. Our results on all the LineMOD objects are detailed in Table S5 of the supplementary material.
@@ IV-D Experimental Results on the LM-O Dataset
Method Settings Error↓ Acc $t^{\circ}$ (%) ↑
@@ V Ablation Analysis
Metrics Mean Err↓ Acc $30^{\circ}$ Acc $15^{\circ}$ Acc $10^{\circ}$ Acc $5^{\circ}$
@@
TABLE VI: Ablation analysis of different semantic features on the LineMOD dataset.
Metrics Mean Err↓ Acc $30^{\circ}$ Acc $15^{\circ}$ Acc $10^{\circ}$ Acc $5^{\circ}$
RGB 36.26 67.42 50.40 37.70 19.62
LoFTR 54.23 46.62 25.00 15.26 5.24
RGB + LoFTR 39.63 64.30 45.14 32.42 14.80
SD 38.45 57.94 37.28 26.32 12.12
RGB + SD 33.78 65.72 47.56 36.76 19.90
DINOv2 31.31 69.32 50.86 38.80 19.22
RGB + DINOv2 29.93 72.06 54.90 42.74 24.32
Figure 8: Illustration of different semantic features as texture maps.
We carefully investigate the following issues by ablation: i) the contribution of each comprising element of our method, including the back-surface culling, and the usage of RGB or semantic modality in Sect. V-A; ii) the ablations on different semantic features in Sect. V-B; iii) the effects of different initialization strategies in Sect. V-C; iv) the effects of different refinement iterations in Sect. V-D; v) the inference time statistics of our method and comparison with other baselines in Sect. V-E; and vi) the failure cases illustrations from the LM-O dataset in Sect. V-F.
V-A The Contributions of the Proposed Comprising Elements
As expected, removing each of our comprising elements results in a decreased performance, because all of them are exploited with clear motivations. Nonetheless, the encouraging observation is that our method is able to deliver promising results using only the RGB modality without the semantic map, which already outperforms the state-of-the-art DVMNet (in-dataset) zhao2024dvmnet in Table II across the rigorous Acc $5^{\circ},10^{\circ},15^{\circ}$, and $30^{\circ}$. This further illustrates the good applicability of our method when the pretrained DINOv2 model is not available.
V-B Ablation on Semantic features
To further investigate the performance incorporating alternative semantic feature representations, we tested semantic features from LoFTR loftr and Stable Diffusion (SD) stablediffusion . Table VI shows that i) LoFTR and SD features are inferior to DINOv2 for this task; ii) RGB-only performance surpasses the results based solely on LoFTR or SD features; iii) complementing LoFTR, SD, or DINOv2 with RGB improves the final performance.
Table VI reveals a significant performance gap among DINOv2, SD, and LoFTR features. To further investigate this, we visualize these features in Fig. 8, where the last three columns are dimensionality-reduced texture maps using PCA. Figure 8 shows that the DINOv2 feature maps best characterize the semantic cues, well complementing the RGB appearance, thus explaining its superior performance.
V-C Effects of Different Initialization Strategies
The pose estimation performance under the render-and-compare paradigm is largely affected by the initialization deepim ; cir ; rnnpose ; megapose ; templates-pose ; latentfusion . In the following, we investigate different initializations including: i) random initialization, where we randomly sample candidate poses and choose the best one; and ii) uniform initialization, where the candidate poses are uniformly sampled from a Fibonacci lattice with in-plane rotations gonzalez2010measurement , as detailed in Sect. III-D (Initialization). Table VII shows the performance of different initialization strategies using the LineMOD dataset, which demonstrates that the uniform initialization outperforms the random initialization.
Moreover, we also perform extensive ablations on the sampling densities using uniform initialization. As shown in Fig. 9, the performance boosts from 0 to 4000 samples, marginally improves from 4000 to 16000, and saturates after 16000 samples. In our experiments, we choose uniform initialization with 4000 samples to balance the performance and the efficiency.
TABLE VII: Effects of different initialization strategies using the LineMOD dataset.
Initial Strategy Sampling Numbers Error↓ Acc @ $t^{\circ}$ (%) ↑
@@ V-C Effects of Different Initialization Strategies
Uniform Init. 4000 29.93 72.06 54.90 42.74 24.32
Figure 9: Ablation analysis on different number of uniformly initialized samples. We tested sampling numbers of 500, 1000, 2000, 4000, 10000, 16000, 32000 on Mean Error and Acc @ $10^{\circ}$.
Figure 10: Performance trend w.r.t. number of refining iterations. We tested iteration numbers of (5, 10, 20, 30, 50, 80, 100, 120, 150) on Mean Error and Acc @ $10^{\circ}$.
V-D Effects of Different Refinement Iterations.
Figure 10 illustrates the impact of the iteration numbers for our label/training-free refinement using the LineMOD dataset, which demonstrates that the performance boosts from 0 to 30 iterations, marginally improves from 30 to 120, and saturates after 120 iterations. We thus set the iteration number to 30 in our main experiments to achieve a balance between performance and efficiency.
@@
0.11s 1.18s 1.02s 2.52s
TABLE IX: Inference time comparison on LineMOD.
Method ZSP RelPose++ LoFTR 3DAHV DMVNet Ours
Time 1.72s 0.69s 0.30s 0.04s 0.04s 2.52s
V-E Analysis on the Inference Time
We collect the inference time per reference-query pair, averaged across the LineMOD datasets on a single 4090 GPU. We report the runtime for each stage of our method in Table VIII. Note that the initialization is efficient with much more candidate samples than the refinement, because those initializing candidate samples can be evaluated in parallel without backpropagation. Table VIII demonstrates the efficiency of our method with a per-pair runtime of 2.52 seconds in total.
We also present comparisons of our approach with the baseline methods, in terms of the inference time, in Table IX. While our method is slower in inference than the state-of-the-art feedforward models, our render-and-compare paradigm is training-free. This eliminates the numerous training hours required by the state-of-the-art feedforward methods, and ensures inherent generalization to unseen objects (i.e., training-free brings desirable training-data independence).
V-F Illustrations of the Failure Cases